Suppose the debt to equity ratio is 2.5 Risk free rate is 3% Equity risk premium is 7%, Beta is 2.0, Cost of debt capital is 5.5%, Effective corporate tax rate is 29%, What is the after-tax WACC?
Answer #1
Source Weightage (A) Post tax cost of capital (B) (A) x (B)
Debt 71.43% 3.91% 71.43% * 3.91% = 2.79%
Equity 28.57% 17.00% 28.57% * 17.00% = 4.86%
WACC = 2.79% + 4.86%
WACC = 7.65%
D/E = 2.5
If Equity = 1
Debt = 1*2.5
Debt = 2.5
Debt to Total capital = 2.5/(2.5+1)
Debt to Total capital = 71.43%
Equity to Total capital = (1 - Debt to Total capital)
Equity to Total capital = (1 - 71.43%)
Equity to Total capital = 28.57%
Cost of equity = Risk free rate + Beta*Equity risk premium
Cost of equity = 3%+(2*7%)
Cost of equity = 17.00%
Post tax cost of debt = Pre-tax cost of debt * (1 - tax rate)
Post tax cost of debt = 5.5%*(1-29%)
Post tax cost of debt = 3.91%